Cohen Company produces and sells socks. Variable cost is $8.50 per pair, and fixed costs for the year total $136,000. The selling price is $17 per pair.

1. Calculate the units required to make a before-tax profit of $76,500. (Do not round intermediate calculations.)

2. Calculate the sales dollars required to make a before-tax profit of $64,600. (Do not round intermediate calculations.)

3. Calculate the sales, in units and in dollars, required to make an after-tax profit of $54,600 given a tax rate of 30%. (Do not round intermediate calculations. Round sales in units up to the nearest whole number and sales in dollars to the nearest whole dollar.)
Answer #1

1. Units required to make a before-tax profit of 76500 $ = Fixed cost + Required before-tax profit / (Selling price - variable cost)

= (136000+76500)/(17-8.5)

= 25000 units

2. Units required to make a before-tax profit of 64600 $ = Fixed cost + Required before-tax profit / (Selling price - variable cost)

= (136000+64600)/(17-8.5)

= 23600 units

3. After tax profit = 54600

Before tax profit = 54600/0.7 = 78000 $

= (136000+78000)/8.5

= 25176.47

= 25176 units